OpenStudy (anonymous):

f(x)=5xsinxcosx, f'(x)= f(x)=1sinxcosx f'(x)=

OpenStudy (experimentx):

do you know product rule?

OpenStudy (anonymous):

yes

OpenStudy (experimentx):

okay let's use product rule for 3 variables. (f(x) g(x) h(x))' = f'(x) g(x) h(x) + f(x) g'(x) h(x) + f(x) g(x) h'(x)

OpenStudy (anonymous):

right,

OpenStudy (anonymous):

so just plug the numbers into the equation

OpenStudy (experimentx):

no ... use f(x) = 5x , g(x) = cos(x) and h(x) = sin(x) on the above formula

OpenStudy (anonymous):

ok, (5x)'(sinx)(cosx)+(5x)(sinx)`(cosx)+(5x)(sinx)(cosx)' ?

OpenStudy (anonymous):

I got (5)(cosx)(sinx)+(5x)(cosx)(cosx)+(5x)(sinx)(sinx)

OpenStudy (experimentx):

this should be +(5x)(sinx)(sinx) = -(5x)(sinx)(sinx)

OpenStudy (anonymous):

ops thank you.

OpenStudy (anonymous):

for the next one how do i do?

OpenStudy (experimentx):

do same for next one ... just use f(x) = 5

OpenStudy (experimentx):

sorry .. 1

OpenStudy (anonymous):

ok, i will wrok on it

OpenStudy (anonymous):

i got (sinx)(cosx)+(cosx^2)-(sinx)^2

OpenStudy (experimentx):

No ... it should be like this 1 sinx cosx = cos^2(x) - sin^2(x)

OpenStudy (anonymous):

how did you get this?

OpenStudy (experimentx):

you didn't use the formula i gave you correctly ... did you?

OpenStudy (anonymous):

ops, you are right, i have another questions

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} tanx/2x\] sin3x/5x

OpenStudy (anonymous):

it says Evaluate the limit

OpenStudy (anonymous):

sec^2/2x?

OpenStudy (experimentx):

do you know that lim x-> 0 sin(x)/x = 1 ?

OpenStudy (anonymous):

Yes,

OpenStudy (experimentx):

write tan(x) as sin(x)/cos(x) ... and then separate it as above.

OpenStudy (anonymous):

sinx/cosx*2x?

OpenStudy (experimentx):

separate the limits like this  Now evaluate the limit individually and multiply them.